Thursday, 9 am MT, 11 am ET, 29 January 2015

Logistics

CV-331/AOC-317/GB-bsm, 331 Hub 192.33.117.12##6519 Audio 434-817-6286

Agenda

  • 15B
    • Helpdesk Tickets (All)
    • PST Review Items (Dana, Jen, Rick)
    • PHT (Barry/Gustaaf)
    • SRP/TAC Recruitment (Mark/Dave/Tim)
    • Science Review Docs (Dana)

  • External/Mixed Joint Proposals (Dana)

  • AOB

Minutes

Attending: BarryClark, RickLively, JenPlank, MarkClaussen, GustaafVanMoorsel, ClaireChandler, DanaBalser, ToneyMInter, DavidFrayer, GarethHunt

  • 15B
    • Helpdesk Tickets. Three issues were discussed.
      • Some issues generating PDFs.
        • A user can save a VLA resource without the frequencies listed. This crashes the PDF reports. Currently you are able to save a VLA resource without the frequencies (or the GOST) image. The frequencies are not trapped at validation, unlike the GOST image. If you then edit this resource again, however, it will check that the frequencies are present when you click save. RickLively will fix this so that the frequencies must be listed when saving during the first edit.
        • If you copy a proposal the GOST image is no longer saved. If you then print the proposal it crashes the PDF reports. JenPlank will try to fix this by either allowing the proposal to be printed or letting the user know with a dialog box (looks like the latter). Regardless, the user cannot submit the proposal since we trap the GOST image at validation.
      • A couple of users have had problems logging into the helpdesk. GarethHunt is following up with these users.
      • There is an issue with the VLA exposure calculator when near the confusion limit and attempting to calculate the time from an input sensitivity. It works the other way around. MarkClaussen is looking into this but it will probably not get fixed soon.
    • PST Review Items. These are almost complete. We agreed to use the following for the default answer to all of the cues in the technical justification review: "All information provided is satisfactory".
    • PHT. The plan is to start working on these items and do as many as possible. It would be useful to have an estimate of how many can be finished. DanaBalser will ping StephanWitz.
    • SRP/TAC Recruitment. MarkClaussen has filled the ETP position with an NRAO person. DavidFrayer is still working to fill the AGN position. He does not want to use an NRAO person since this panel already includes two (one is for 15B only).
    • Science Review Docs. These are now in Plone at: https://science.nrao.edu/observing/proposal-types/documentation. TimBastian has reviewed these. Please have a look, especially at the Technical review document. (These have now been published.)

  • External/Mixed Joint Proposals. An email exploder has been setup. The relevant people at HST, Swift, Chandra, and Fermi have been contacted to use this email address. They have all agreed to give us 3-4 weeks for performing technical reviews on external joint proposals, and they have agreed to complete technical reviews for internal joint proposals by the time of the SRP telecon.

  • AOB
    • Please forward any absences.

-- DanaBalser - 2015-01-28
Topic revision: r2 - 2015-01-29, DanaBalser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback